# Robust Preference Optimization: Aligning Language Models with Noisy Preference Feedback

### v1

@@ Abstract
Standard human preference-based alignment methods, such as Reinforcement Learning from Human Feedback (RLHF), are a cornerstone technology for aligning Large Language Models (LLMs) with human values. However, these methods are all underpinned by a critical, yet flawed assumption: human preferences are homogeneous (representing a single, unified preference) and the collected data is noiseless (free from error). In reality, neither is true since human preference is pluralistic and annotators can make mistakes. This creates a discrepancy between the recorded data and the ground-truth preferences, which can misguide the model and degrade its performance. To address this challenge, we introduce Robust Preference Optimization (RPO). RPO employs an Expectation-Maximization (EM) algorithm to infer the posterior probability of each label’s correctness, which is used to adaptively re-weigh each data point in the training loss to mitigate noise. We further generalize this approach by establishing a theoretical link between arbitrary preference losses and their corresponding probabilistic models. This generalization enables the systematic transformation of existing alignment algorithms into their robust counterparts, elevating RPO from a specific algorithm to a meta-framework for robust preference alignment. Theoretically, we prove that under the condition of a perfectly calibrated model, RPO is guaranteed to converge to the true noise level of the dataset. Our experiments demonstrate RPO’s effectiveness as a meta-framework, consistently enhancing four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO). When applied to Mistral and Llama 3 models, the RPO-enhanced methods achieve substantial win rate gains on AlpacaEval 2 and Arena-Hard, with improvements of up to 7.0% and 5.4%, respectively.
@@ 1 Introduction
Figure 1: The Robust Preference Optimization (RPO) framework iteratively learns a collective preference from noisy data using an Expectation-Maximization (EM) algorithm. In each cycle, the E-step computes a confidence score for each label by inferring the posterior probability of the true underlying preference. The M-step then uses these scores as adaptive weights to update both the LLM policy and the annotator reliability parameters. This process repeats, allowing the policy and reliability scores to mutually improve.
@@
To address this challenge, we propose Robust Preference Optimization (RPO). Instead of treating labels as a fixed truth, our approach aims to learn a robust model of collective preference that best explains the entire noisy and diverse dataset. The core innovation of RPO is its departure from the hard labels used in traditional RLHF. We treat the true preference as a latent variable and compute soft preference labels to estimate the underlying consensus. To achieve this, RPO employs an Expectation-Maximization (EM) framework that simultaneously models annotator reliability while optimizing the LLM. In the E-step, it infers the posterior probability that each annotated label is correct, effectively estimating annotator reliability. In the M-step, it uses these probabilities as adaptive weights to update the LLM, thereby learning from a dynamically re-weighted preference signal.
Our comprehensive experiments validate RPO as a highly effective and versatile meta-framework. We show that applying RPO consistently enhances four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO) across two different base models (Mistral-7B and Llama-3-8B) on two widely-used benchmarks (Table 2). RPO-enhanced methods achieve substantial win-rate gains over their standard counterparts, with improvements of up to 7.0% on AlpacaEval 2 and 5.4% on Arena-Hard. Furthermore, we theoretically prove that RPO can recover the true reliability of annotators (Theorem 4.2) and empirically verify this guarantee in controlled experiments (section 5.4).
@@
  • We propose Robust Preference Optimization, an algorithm that achieves robust LLM alignment by actively denoising preference data. RPO formulates this as a latent variable problem, employing an EM approach to infer the reliability of each label while simultaneously optimizing the LLM policy.
  • We theoretically establish a generalized RPO framework by using the Gibbs distribution to connect any preference loss function to an underlying probabilistic model. This generalization elevates RPO to a meta-framework, capable of systematically transforming existing alignment methods into their robust counterparts.
  • We conduct extensive experiments that demonstrate the practical effectiveness and versatility of our framework. By applying RPO to several popular alignment methods, we show significant performance improvements, providing a unified approach to robust LLM alignment.
@@ 2 Related work
The standard paradigm for aligning Large Language models (LLMs) with human values is Reinforcement Learning from Human Feedback (RLHF), which involves training a reward model and then fine-tuning the policy against it (Christiano et al., 2017; Ouyang et al., 2022). To mitigate the complexity and instability of this multi-stage process, a family of simpler, direct alignment algorithms has emerged (Rafailov et al., 2023; Azar et al., 2023; Meng et al., 2024; Hong et al., 2024). These methods bypass the explicit reward modeling stage by optimizing a direct classification-style loss on the preference data. However, a critical limitation shared by these methods is their reliance on hard preference labels. This approach models human feedback as a definitive, binary choice, treating every label with equal and absolute confidence. Consequently, it is highly vulnerable to the significant label noise present in real-world datasets, as standard loss functions can lead models to overfit to corrupted labels (Natarajan et al., 2013; Zhang and Sabuncu, 2018). A simple annotation error, such as an accidental misclick, is given the same weight as a deliberate, high-quality judgment. This inability to distinguish between reliable feedback and noise means that the model’s performance degrades significantly as the error rate increases (Frénay and Verleysen, 2013; Gao et al., 2024). In contrast, a soft preference label approach, which models preferences probabilistically, can better handle this uncertainty. By assigning a confidence score or weight to each label, the model can learn to rely more on high-quality signals while down-weighting probable noise (Müller et al., 2019; Song et al., 2024). This is the core advantage of our work; RPO outperforms prior methods because its soft-label framework explicitly models and mitigates the impact of label noise, leading to a more robust and effective alignment.
@@
The vulnerability to label noise situates preference alignment within the classic machine learning problem of Learning with Noisy Labels (LNL) (Natarajan et al., 2013). Foundational work in LNL, such as the Dawid and Skene model (Dawid and Skene, 1979), used an EM algorithm to simultaneously infer true labels while estimating annotator reliability. This principle was later extended to pairwise comparisons in the Crowd-BT model (Chen et al., 2013), which jointly estimates item scores and annotator reliability in crowdsourced ranking tasks. In the context of modern LLM alignment, several approaches have recently been proposed to address label noise. These can be broadly categorized into methods that design robust loss functions and those that implement data-centric filtering strategies. In the first category, rDPO (Chowdhury et al., 2024) constructs an unbiased estimator of the true loss but requires the global noise rate to be known a priori. Hölder-DPO (Fujisawa et al., 2025) introduces a loss with a ”redescending property,” which inherently nullifies the influence of extreme outliers without needing a known noise rate. In the second category, Selective DPO (Gao et al., 2025) proposes filtering examples based on their difficulty relative to the model’s capacity—a concept orthogonal to label correctness—using validation loss as a proxy. Another recent framework, ORPO (Hong et al., 2024), employs a hybrid iterative approach that alternates between training with a robust loss and explicitly filtering out a fixed fraction of high-loss samples. Our proposed RPO framework is distinct from these contemporary methods. Instead of merely mitigating the consequences of noise (e.g., high-loss samples), RPO addresses its root cause by directly modeling the data-generating process. By treating annotator reliability as a latent variable to be inferred, it can distinguish signal from noise in a principled manner, offering a more robust path to alignment.
@@ 3 Methodology
This section details our proposed RPO algorithm. We begin by briefly reviewing the standard DPO framework, which serves as our baseline in section 3.1. We then identify its key limitation in handling noisy data, which motivates the core of our work: a new framework that treats preferences as latent variables. Finally, we derive the RPO algorithm using an Expectation-Maximization (EM) approach and discuss its practical implementation.
@@ 3.1 Preliminaries: Direct Preference Optimization
$\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
@@
3.2 RPO Framework: Core Assumptions
A critical limitation of DPO is its implicit assumption that all observed preferences in $\mathcal{D}$ are correct. In practice, this data is often noisy. To address this, we propose Robust Preference Optimization (RPO), which is built upon two core assumptions that reframe the problem.
@@ 3.2 RPO Framework: Core Assumptions
4      For each sample $i\in\mathcal{B}$, compute $w_{i}$ using current $\theta$ and $\eta_{k_{i}}$ via (4);
@@
7     Compute the weighted loss $\mathcal{L}_{\text{RPO}}(\theta)$ for the batch via (5);
@@ 11         Update $\eta_{k}$ via (7);
Algorithm 1 Robust Preference Optimization (RPO)
@@
3.3 The RPO Algorithm via Expectation-Maximization
@@
$w_{i}^{(t)}\leftarrow\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})}.$ (4)
where $p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})$ and $p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})$ can be computed according to (2).
@@
First, the policy is updated by minimizing a weighted loss function. As established in Assumption 2, our probabilistic model for $p(y_{w}\succ^{\ast}y_{l})$ allows RPO to work with any preference loss $\mathcal{L}_{\text{pref}}$, making it a versatile meta-framework. The general RPO loss is:
$\min_{\theta}\mathcal{L}_{\text{RPO}}(\theta)=-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)\right]$ (5)
@@
$\eta_{k}^{(t+1)}\leftarrow\frac{\sum_{i\in\mathcal{I}_{k}}w_{i}^{(t)}}{N_{k}}$ (6)
@@ 3.4 Practical implementation with mini-batch training
$\eta_{k}\leftarrow(1-\alpha)\eta_{k}+\alpha\cdot\frac{\sum_{i\in\mathcal{B}\cap\mathcal{I}_{k}}w_{i}}{N_{k,\mathcal{B}}}$ (7)
Here, $N_{k,\mathcal{B}}$ is the number of examples from annotator $k$ in the current mini-batch, and $\alpha\in(0,1]$ is a momentum hyperparameter. The complete training procedure for RPO is summarized in Algorithm 1:
4 Theoretical analysis of RPO
The robustness of RPO stems from its adaptive weighting mechanism. This section first provides an intuitive analysis of these training dynamics and then formalizes this intuition with theoretical guarantees, demonstrating that the RPO framework can recover the true reliability of annotators.
@@ 4 Theoretical analysis of RPO
We now formalize the intuition that RPO can recover the true reliability of annotators. We provide this analysis under an idealized setting: full-batch training where the M-step for the policy parameters $\theta$ is assumed to have converged perfectly. While our practical implementation in Algorithm 1 uses mini-batch gradient updates (a form of Generalized EM), this idealized analysis provides a strong theoretical justification for our framework.
@@
Theorem 4.2 (Identification and convergence of RPO).
@@
5 Experiments
Table 2: Performance comparison on AlpacaEval 2 (Dubois et al., 2024) and Arena-Hard (Li et al., 2024). LC and WR denote length-controlled and raw win rate, respectively. Each pair consists of a standard alignment algorithm and its robust counterpart using our RPO framework (e.g., R-DPO). We also report the performance of the base models without any fine-tuning for reference. Our RPO-enhanced methods consistently outperform the base algorithms. The best scores in each pairing are in bold, with absolute gains from RPO highlighted in red.
@@ 5 Experiments
R-CPO (Ours) 27.6 (+1.3) 27.8 (+1.4) 14.7 (+1.2) 40.1 (+4.2) 43.8 (+3.5) 28.3 (+2.6)
In this section, we conduct a comprehensive set of experiments to evaluate the preformance of RPO. We begin in section 5.1 by detailing our experimental setup, including the models, datasets, evaluation benchmarks, and baseline algorithms. In Section 5.2, we present our main results. We then conduct an ablation study in Section 5.3 to analyze the framework’s sensitivity to its key hyperparameters. In section 5.4, we provide an empirical verification of our theoretical claims from Theorem 4.2.

### v2

# Latent Collective Preference Optimization: A General Framework for Robust LLM Alignment

@@ Abstract
Standard human preference-based alignment methods, such as Reinforcement Learning from Human Feedback (RLHF), are a cornerstone technology for aligning Large Language Models (LLMs) with human values. However, these methods are all underpinned by a critical, yet flawed assumption: human preferences are homogeneous (representing a single, unified preference) and the collected data is noiseless (free from error). In reality, neither is true since human preference is pluralistic and annotators can make mistakes. This creates a discrepancy between the recorded data and the ground-truth preferences, which can misguide the model and degrade its performance. To address this challenge, we introduce Latent Collective Preference Optimization (LCPO). LCPO leverages an Expectation-Maximization (EM) algorithm to learn the latent collective consensus from noisy data. It operates by inferring the correctness of each preference label and using this probability as an adaptive weight to re-calibrate each data point’s contribution to the training loss, thereby mitigating noise. We generalize this approach by establishing a theoretical link between arbitrary preference losses and their corresponding probabilistic models, elevating LCPO from a specific algorithm to a general framework for robust preference alignment. Theoretically, we prove that under the condition of a perfectly calibrated model, LCPO is guaranteed to converge to the true noise level of the dataset. Our experiments demonstrate LCPO’s effectiveness as a general framework, consistently enhancing four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO). When applied to Mistral and Llama 3 models, the LCPO-enhanced methods achieve substantial win rate gains on AlpacaEval 2 and Arena-Hard, with improvements of up to 7.0% on both benchmarks.
@@ 1 Introduction
Figure 1: The Latent Collective Preference Optimization (LCPO) framework iteratively learns a collective preference from noisy data using an Expectation-Maximization (EM) algorithm. In each cycle, the E-step computes a confidence score for each label by inferring the posterior probability of the true underlying preference. The M-step then uses these scores as adaptive weights to update both the LLM policy and the annotator reliability parameters. This process repeats, allowing the policy and reliability scores to mutually improve.
@@
To address this challenge, we propose Latent Collective Preference Optimization (LCPO). Instead of treating labels as a fixed truth, our approach aims to learn a robust model of collective preference that best explains the entire noisy and diverse dataset. The core innovation of LCPO is its departure from the hard labels used in traditional RLHF. We treat the true preference as a latent variable and compute soft preference labels to estimate the underlying consensus. To achieve this, LCPO employs an Expectation-Maximization (EM) framework that simultaneously models annotator reliability while optimizing the LLM. In the E-step, it infers the posterior probability that each annotated label is correct, effectively estimating annotator reliability. In the M-step, it uses these probabilities as adaptive weights to update the LLM, thereby learning from a dynamically re-weighted preference signal.
Our comprehensive experiments validate LCPO as a highly effective and versatile meta-framework. We show that applying LCPO consistently enhances four state-of-the-art alignment algorithms (DPO, IPO, SimPO, and CPO) across two different base models (Mistral-7B and Llama-3-8B) on two widely-used benchmarks (Table 2). LCPO-enhanced methods achieve substantial win-rate gains over their standard counterparts, with improvements of up to 7.0% on AlpacaEval 2 and 5.4% on Arena-Hard. Furthermore, we theoretically prove that LCPO can recover the true reliability of annotators (Theorem 4.2) and empirically verify this guarantee in controlled experiments (Section 5.4).
@@
  • We propose Latent Collective Preference Optimization, an algorithm that achieves robust LLM alignment by actively denoising preference data. LCPO formulates this as a latent variable problem, employing an EM approach to infer the reliability of each label while simultaneously optimizing the LLM policy.
  • We theoretically establish a generalized LCPO framework by using the Gibbs distribution to connect any preference loss function to an underlying probabilistic model. This generalization elevates LCPO to a meta-framework, capable of systematically transforming existing alignment methods into their robust counterparts.
  • We conduct extensive experiments that demonstrate the practical effectiveness and versatility of our framework. By applying LCPO to several popular alignment methods, we show significant performance improvements, providing a unified approach to robust LLM alignment.
@@ 2 Related work
The standard paradigm for aligning Large Language models (LLMs) with human values is Reinforcement Learning from Human Feedback (RLHF), which involves training a reward model and then fine-tuning the policy against it (Christiano et al., 2017; Ouyang et al., 2022). To mitigate the complexity and instability of this multi-stage process, a family of simpler, direct alignment algorithms has emerged (Rafailov et al., 2023; Azar et al., 2023; Meng et al., 2024; Hong et al., 2024). These methods bypass the explicit reward modeling stage by optimizing a direct classification-style loss on the preference data. However, a critical limitation shared by these methods is their reliance on hard preference labels. This approach models human feedback as a definitive, binary choice, treating every label with equal and absolute confidence. Consequently, it is highly vulnerable to the significant label noise present in real-world datasets, as standard loss functions can lead models to overfit to corrupted labels (Natarajan et al., 2013; Zhang and Sabuncu, 2018). A simple annotation error, such as an accidental misclick, is given the same weight as a deliberate, high-quality judgment. This inability to distinguish between reliable feedback and noise means that the model’s performance degrades significantly as the error rate increases (Frénay and Verleysen, 2013; Gao et al., 2024). In contrast, a soft preference label approach, which models preferences probabilistically, can better handle this uncertainty. By assigning a confidence score or weight to each label, the model can learn to rely more on high-quality signals while down-weighting probable noise (Müller et al., 2019; Song et al., 2024). This is the core advantage of our work; LCPO outperforms prior methods because its soft-label framework explicitly models and mitigates the impact of label noise, leading to a more robust and effective alignment.
@@
The vulnerability to label noise situates preference alignment within the classic machine learning problem of Learning with Noisy Labels (LNL) (Natarajan et al., 2013). Foundational work in LNL, such as the Dawid and Skene model (Dawid and Skene, 1979), used an EM algorithm to simultaneously infer true labels while estimating annotator reliability. This principle was later extended to pairwise comparisons in the Crowd-BT model (Chen et al., 2013), which jointly estimates item scores and annotator reliability in crowdsourced ranking tasks. In the context of modern LLM alignment, several approaches have recently been proposed to address label noise. These can be broadly categorized into methods that design robust loss functions and those that implement data-centric filtering strategies. In the first category, rDPO (Chowdhury et al., 2024) constructs an unbiased estimator of the true loss but requires the global noise rate to be known a priori. Hölder-DPO (Fujisawa et al., 2025) introduces a loss with a ”redescending property,” which inherently nullifies the influence of extreme outliers without needing a known noise rate. In the second category, Selective DPO (Gao et al., 2025) proposes filtering examples based on their difficulty relative to the model’s capacity—a concept orthogonal to label correctness—using validation loss as a proxy. Another recent framework, ORPO (Hong et al., 2024), employs a hybrid iterative approach that alternates between training with a robust loss and explicitly filtering out a fixed fraction of high-loss samples. ROPO (Liang et al., 2024) proposes an iterative framework that also alternates between noise-tolerant training and noisy sample filtering. Our proposed LCPO framework is distinct from these contemporary methods. Instead of merely mitigating the consequences of noise (e.g., high-loss samples), LCPO addresses its root cause by directly modeling the data-generating process. By treating annotator reliability as a latent variable to be inferred, it can distinguish signal from noise in a principled manner, offering a more robust path to alignment.
@@ 3 Methodology
This Section details our proposed LCPO algorithm. We begin by briefly reviewing the standard DPO framework, which serves as our baseline in Section 3.1. We then identify its key limitation in handling noisy data, which motivates the core of our work: a new framework that treats preferences as latent variables. Finally, we derive the LCPO algorithm using an Expectation-Maximization (EM) approach and discuss its practical implementation.
@@ 3.1 Preliminaries: Direct Preference Optimization
$\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (1)
@@
3.2 LCPO Framework: Core Assumptions
A critical limitation of DPO is its implicit assumption that all observed preferences in $\mathcal{D}$ are correct. In practice, this data is often noisy. To address this, we propose Latent Collective Preference Optimization (LCPO), which is built upon two core assumptions that reframe the problem.
@@ 3.2 LCPO Framework: Core Assumptions
4      For each sample $i\in\mathcal{B}$, compute $w_{i}$ using current $\theta$ and $\eta_{k_{i}}$ via equation (4);
@@
7     Compute the weighted loss $\mathcal{L}_{\text{LCPO}}(\theta)$ for the batch via equation (5);
@@ 11         Update $\eta_{k}$ via equation (7);
Algorithm 1 Latent Collective Preference Optimization (LCPO)
@@
3.3 The LCPO Algorithm via Expectation-Maximization
@@
$w_{i}^{(t)}\leftarrow\frac{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}}{p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})\eta_{k_{i}}^{(t)}+p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})(1-\eta_{k_{i}}^{(t)})},$ (4)
where $p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta^{(t)})$ and $p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta^{(t)})$ can be computed according to equation (2).
@@
First, the policy is updated by minimizing a weighted loss function. As established in Assumption 2, our probabilistic model for $p(y_{w}\succ^{\ast}y_{l})$ allows LCPO to work with any preference loss $\mathcal{L}_{\text{pref}}$, making it a versatile meta-framework. The general LCPO loss is:
$\min_{\theta}\mathcal{L}_{\text{LCPO}}(\theta)=-\sum_{i=1}^{N}\left[w_{i}^{(t)}\log p(y_{w,i}\succ^{\ast}y_{l,i}|x_{i},\theta)+(1-w_{i}^{(t)})\log p(y_{l,i}\succ^{\ast}y_{w,i}|x_{i},\theta)\right].$ (5)
@@
$\eta_{k}^{(t+1)}\leftarrow\frac{\sum_{i\in\mathcal{I}_{k}}w_{i}^{(t)}}{N_{k}}.$ (6)
@@ 3.4 Practical implementation with mini-batch training
$\eta_{k}\leftarrow(1-\alpha)\eta_{k}+\alpha\cdot\frac{\sum_{i\in\mathcal{B}\cap\mathcal{I}_{k}}w_{i}}{N_{k,\mathcal{B}}},$ (7)
Here, $N_{k,\mathcal{B}}$ is the number of examples from annotator $k$ in the current mini-batch, and $\alpha\in(0,1]$ is a momentum hyperparameter. The complete training procedure for LCPO is summarized in Algorithm 1:
4 Theoretical analysis of LCPO
The robustness of LCPO stems from its adaptive weighting mechanism. This section first provides an intuitive analysis of these training dynamics and then formalizes this intuition with theoretical guarantees, demonstrating that the LCPO framework can recover the true reliability of annotators.
@@ 4 Theoretical analysis of LCPO
We now formalize the intuition that LCPO can recover the true reliability of annotators. We provide this analysis under an idealized setting: full-batch training where the M-step for the policy parameters $\theta$ is assumed to have converged perfectly. While our practical implementation in Algorithm 1 uses mini-batch gradient updates (a form of Generalized EM), this idealized analysis provides a strong theoretical justification for our framework.
@@
Theorem 4.2 (Identification and convergence of LCPO).
@@
Table 2: Performance comparison on AlpacaEval 2 (Dubois et al., 2024) and Arena-Hard (Li et al., 2024). LC and WR denote length-controlled and raw win rate, respectively. Each pair consists of a standard alignment algorithm and its robust counterpart using our LCPO framework (e.g., LCPO-DPO). We also report the performance of the base models without any fine-tuning for reference. Our LCPO-enhanced methods consistently outperform the base algorithms. The best scores in each pairing are in bold, with absolute gains from LCPO highlighted in red.
@@
LCPO-CPO (Ours) 27.6 (+1.3) 27.8 (+1.4) 14.7 (+1.2) 40.1 (+4.2) 43.8 (+3.5) 28.3 (+2.6)
5 Experiments
In this section, we conduct a comprehensive set of experiments to evaluate the preformance of LCPO. We begin in Section 5.1 by detailing our experimental setup, including the models, datasets, evaluation benchmarks, and baseline algorithms. In Section 5.2, we present our main results. We then conduct an ablation study in Section 5.3 to analyze the framework’s sensitivity to its key hyperparameters. In Section 5.4, we provide an empirical verification of our theoretical claims from Theorem 4.2.